OpenStudy (shamil98):

∫ |x| = |x^2/2| + c?

OpenStudy (anonymous):

integrate(|x|)=x^3/(2|x|)+c

OpenStudy (anonymous):

this can be done piecewise: 1) x>0: integrate(|x|)=integrate(x)=x^2/2 2) x<0: integrate(|x|)=integrate(-x)=-x^2/2 3) integrate(|x|)=x^3/(2|x|)+c

OpenStudy (shamil98):

I don't get the third method... why would it be x^3/2(|x|) + c ?

OpenStudy (anonymous):

if x>0, then x^3/(2|x|) equals x^2/2 if x<0, then x^3/(2|x|) equals -x^2/2

OpenStudy (anonymous):

the answer could also be written as |x^3|/(2x)+c

OpenStudy (shamil98):

Okay, if I had the bounds of (-2, 1) and the equation was: ∫ |x| dx , I would use the fundamental theorem of calculus then right?

OpenStudy (anonymous):

yup. either that, or break the integral up into two parts: integrate(-x from x=-2 to x=0)+integrate(x from x=0 to x=1)

OpenStudy (shamil98):

alright, thanks a lot!